Part A Part complete What is the magnitude of the electric field? E = 559 V/m...

Part A

Part complete

What is the magnitude of the electric field?

E =

559

V/m

SubmitPrevious Answers

Correct

Part B

Part complete

What is the direction of the electric field?

θ =

243

∘ from the +x axis

SubmitPrevious Answers

Correct

Part C

What is the shortest distance one can move to undergo a change in potential of 2.00 V ?

d

d

=

SubmitPrevious AnswersRequest Answer

Incorrect; Try Again; One attempt remaining
Answer #1

(C) We have the relation:

E = V/d

Rearranging in terms of d, we get:

d = V/E

= 2/559

= 3.58e-3 m

Kindly rate :)